OpenStudy (anonymous):

An airplane leaves an airport and flies due west 120 miles and then 150 miles in the direction S 39.17°W. How far is the plane from the airport (round to the nearest mile)? PLEASE HELP I GIVE MEDALS

OpenStudy (anonymous):

Need the work ... PLease help Im confused

OpenStudy (anonymous):

Im really not sure :( its in my segment about law of cosines

OpenStudy (anonymous):

draw a diagram

OpenStudy (anonymous):

This is the picture one sec

OpenStudy (anonymous):

Here it is @artofspeed

OpenStudy (anonymous):

thats correct

OpenStudy (anonymous):

you gotta use cos law for this one

OpenStudy (anonymous):

@artofspeed can you help me with the work .. I have to answer the question and sow work

OpenStudy (anonymous):

do you know the law of cos

OpenStudy (anonymous):

isnt it c^2 = a^2 +b^2

OpenStudy (anonymous):

no, that's Pythagoras

OpenStudy (anonymous):

Then no :(

OpenStudy (anonymous):

or a2 + b2 - 2ab cos(C) = c2 ?

OpenStudy (anonymous):

yesssssssssssssssss

OpenStudy (anonymous):

here, the angle C is 39.17+90, and u have a, b, only c left to find, which is the answer

OpenStudy (anonymous):

so a = 120 and b = 150 and angle C = 39.17+90

OpenStudy (anonymous):

how do I do that @artofspeed ? ps thank you so far

OpenStudy (anonymous):

replace them in a2 + b2 - 2ab cos(C) = c2 to find c

OpenStudy (anonymous):

and c is the final answer

OpenStudy (anonymous):

so I would have 120^2 + 150^2 -2(120)(150)cos(129.17)=c^2? @artofspeed

OpenStudy (anonymous):

Hello??? ;( @artofspeed

OpenStudy (anonymous):

yes, now find c

OpenStudy (anonymous):

I kept getting something wrong ... Im not sure what to do now to find it ? @artofspeed im doing something wrong ...c an you walk me through it ?

OpenStudy (anonymous):

I think I know the answer is 244 mi but I dont know how to explain it

OpenStudy (anonymous):

do you have the solution for this problem

OpenStudy (anonymous):

I have the answer but my I need the work so I know how I did it and Im not sure how to get c @artofspeed

OpenStudy (anonymous):

wait, did you get the correct answer or not..

OpenStudy (anonymous):

@mathstudent55 anything yet ? I just got back

OpenStudy (mathstudent55):